In Desmos:

write all transformation that occurred in the parent function y=square root of x


Transformed function: y= - square root x-3 +1

1
Expert's answer
2021-03-07T14:00:34-0500

Solution.


"y=\sqrt{x}" (blue line)

"y=-\sqrt{x}" - symmetry about the x-axis (green line)

"y=-\sqrt{x-3}" - parallel translation by 3 units to the right (purple line)

"y=-\sqrt{x-3}+1" - parallel translation by 1 unit up (red line)